OpenStudy (lgbasallote):

\[\textbf{X} - \left[ \begin{array}{ccc} 0 & 1 & 2 \\ 3 & 4 & 5 \\ 6 & 7 &8i \end{array} \right] = \left[ \begin{array}{ccc} 8 & 7 & 6 \\ 5 & 4 & 3 \\ 2 & 1 & 0 \end{array} \right]\] basically i need helpwith operations of matrices...

OpenStudy (unklerhaukus):

just add the first matrix to both sides

OpenStudy (lgbasallote):

uhmm E33 us 8 not 8i

OpenStudy (anonymous):

ahahah (dumudugo utak koh.. ) xD

OpenStudy (lgbasallote):

@UnkleRhaukus like i said what i need help on is how to add matrices

OpenStudy (lgbasallote):

lol @somebody11111 i think this is basic stuff :p

OpenStudy (unklerhaukus):

just add the elements

OpenStudy (lgbasallote):

well for college..

Parth (parthkohli):

Just add the corresponding elements.

OpenStudy (lgbasallote):

uhmm again..i ask... *how*

OpenStudy (anonymous):

out na akoh.. baka buong katawan ko naman dumugo. xD

Parth (parthkohli):

Add 0 + 8 Add 1 + 7 etc and then place them in the corresponding positions in the matrix

OpenStudy (lgbasallote):

if i knew how to add 1 + 1 i wouldnt ask 1 + 1 lol..i'll expect someone to draw sticks to show 1 + 1 = 2

OpenStudy (lgbasallote):

@ParthKohli seriously...HOW

OpenStudy (lgbasallote):

im so tired of repeating the word =_=

Parth (parthkohli):

A1 + A1 ===> Put that A1 in the new matrix just like that

OpenStudy (unklerhaukus):

\[\textbf{X} - \left[ \begin{array}{ccc} 0 & 1 & 2 \\ 3 & 4 & 5 \\ 6 & 7 &8i \end{array} \right] = \left[ \begin{array}{ccc} 8 & 7 & 6 \\ 5 & 4 & 3 \\ 2 & 1 & 0 \end{array} \right]\] \[\textbf{X} = \left[ \begin{array}{ccc} 0 & 1 & 2 \\ 3 & 4 & 5 \\ 6 & 7 &8 \end{array} \right] +\left[ \begin{array}{ccc} 8 & 7 & 6 \\ 5 & 4 & 3 \\ 2 & 1 & 0 \end{array} \right]\] \[\textbf{X} = \left[ \begin{array}{ccc} 0 +8& 1+7 & 2+6 \\ 3+5 & 4+4 & 5+3 \\ 6+2 & 7+1 &8+0 \end{array} \right] \]

Parth (parthkohli):

Did you understand? @lgbasallote

OpenStudy (unklerhaukus):

get it?

OpenStudy (lgbasallote):

ahh so it's just like that..thanks!!

OpenStudy (lgbasallote):

yep!!!

OpenStudy (lgbasallote):

subtraction is just the same right?

Parth (parthkohli):

Yeah

OpenStudy (lgbasallote):

ok thanks again

OpenStudy (unklerhaukus):

note : you can only add matrices of the same size

OpenStudy (lgbasallote):

hmm?

OpenStudy (unklerhaukus):

yep subtraction is just the same

OpenStudy (unklerhaukus):

\[\textbf{X} = \left[ \begin{array}{ccc} 8 & 8 &8 \\ 8 & 8 & 8 \\ 8 & 8 &8 \end{array} \right] \]

OpenStudy (lgbasallote):

so what happens when it's 2 x 2 + 3 x 3? it remains the same?

Parth (parthkohli):

You can't add: \( \color{Black}{\Rightarrow \Large \left [\begin{array}{} {a \\ b} \end{array} \right ] + \left [ \begin{array}{} {c \\ d \\ e}\end{array}\right ]}\) Etc.

OpenStudy (unklerhaukus):

\[\textbf{X} = 8\left[ \begin{array}{ccc} 1 & 1 &1 \\ 1 & 1 & 1 \\ 1 & 1 &1 \end{array} \right]\]

OpenStudy (lgbasallote):

ahh that answers my second question..

OpenStudy (lgbasallote):

constant * nxn matrix

OpenStudy (lgbasallote):

so what happens in 2 x 2 + 3 x 3? it remains the same???

OpenStudy (unklerhaukus):

you can not simplify \[\textbf A_{nm}+\textbf B_{pq}\] if n≠p and/or m≠q

OpenStudy (lgbasallote):

so...it remains the same? like x + y?